Find the integral surface of the PDE:


x²p + y²q +z² =0


which passes through the hyperbola


xy =x+y, z=1.

1
Expert's answer
2022-02-22T08:36:28-0500

The auxiliary equation is "\dfrac{dx}{x^2}=\dfrac{dy}{y^2}=\dfrac{dz}{-z^2}."

Taking the first two ratios,


"\begin{aligned}
\dfrac{dx}{x^2}&=\dfrac{dy}{y^2}\\
\text{Integrating both sides, we get}\\
\dfrac{-1}{x} &= \dfrac{-1}{y} + c_{1}\\
\dfrac{1}{y} -\dfrac{1}{x} &= c_{1}\qquad\qquad(1)\\
\end{aligned}"

Taking last two ratios,


"\begin{aligned}
\dfrac{dy}{y^2}&=-\dfrac{dz}{z^2}\\
\text{Integrating both sides, we get}\\
\dfrac{-1}{y} &= \dfrac{1}{z} + c_{2}\\
-\dfrac{1}{y} -\dfrac{1}{z} &= c_{2}\qquad\qquad(2)\\
\end{aligned}"


The integral curve is "\phi(c_1,c_2) = 0", i.e., "\phi(\frac{1}{y}-\frac{1}{x},-\frac{1}{y}-\frac{1}{z}) = 0".


Given that the integral surface passes through "xy=x+y, z=1".


Letting "x(s) = s" as the parametric representation, we get

"\begin{aligned}
sy &= s+y\\
sy-y&=s\\
y(s-1) &= s\\
\therefore y(s) &= \dfrac{s}{s-1}
\end{aligned}"

Using "x(s) = s, y(s) = \dfrac{s}{s-1}, z=1" in (1) and (2) we get

"\begin{aligned}
\frac{s-1}{s}-\frac{1}{s} &= c_1 \Rightarrow c_1 = \dfrac{s-2}{s}\Rightarrow c_1 = 1- \dfrac{2}{s} \Rightarrow \dfrac{1}{s} = \dfrac{1-c_1}{2}\qquad (3)\\
-\left(\frac{s-1}{s}\right)-1 &= c_2\Rightarrow c_2 = \dfrac{1-2s}{s} \Rightarrow c_2 = \dfrac{1}{s}-2 \qquad\qquad\qquad\qquad(4)\\
\end{aligned}"


Using (3) in (4),

"\begin{aligned}
c_2 &= \dfrac{1-c_1}{2} - 2\\
2c_2 - 1 + c_1+4 &= 0\\
c_1 + 2c_2+3 &=0\\
\left(\dfrac{1}{y}-\dfrac{1}{x}\right) - 2\left(\dfrac{1}{y}+\dfrac{1}{z}\right)+3&=0~~~(\text{Using (1) \& (2)})\\
\dfrac{x-y}{xy}-2\left(\dfrac{z+y}{yz}\right)+3&=0\\
xz-yz-2xz-2xy+3xyz &=0\\
3xyz-2xy-xz-yz&=0
\end{aligned}"

which is the required integral surface.